OpenStudy (anonymous):

Find the linear approximation of f(x)=lnx at x=1 and use it to estimate ln1.38. L(x)= ? ln1.38 approximately = ?

OpenStudy (amistre64):

the equation of the tangent to the curve is the approx

OpenStudy (anonymous):

But I don't know how to do that, can you please help?

OpenStudy (amistre64):

which part, finding a derivative, or working that into a linear equation?

OpenStudy (anonymous):

Linear equation

OpenStudy (amistre64):

y = mx - mPx + Py is a good rule to follow; this is where m=slope, and we know a point (Px,Py)

OpenStudy (amistre64):

the derivative IS the slope of the curve ; y = f'(x)x - f'(x) Px + Py Px is simply the value ofx in the point used, in this case 1 y = f'(1)x -f'(1)(1) + Py and Py is the value of the function at x=1; or simply .. f(1) y = f'(1)x -f'(1) + f(1)

OpenStudy (amistre64):

what is the derivate of ln(x) ?

OpenStudy (anonymous):

1/x

OpenStudy (anonymous):

By y do you mean f(x)?

OpenStudy (anonymous):

I mean, by y do you mean L(x)?

OpenStudy (amistre64):

yes, the linear function is y = L(x)

OpenStudy (amistre64):

1/x at x=1 is 1

OpenStudy (amistre64):

L(x) = f'(1)x -f'(1) + f(1) f'(1) = 1 L(x) = x - 1 + f(1) and f(1) = ln(1) = 0 L(x) = x - 1 + 0 L(x) = x - 1 is the linear equation approximation

OpenStudy (amistre64):

seems a little off .....

OpenStudy (amistre64):

i see it; ln(1) = 0; 1-1=0 ... its good

OpenStudy (amistre64):

ln(1.38) = abt 1.38-1 = .38

OpenStudy (anonymous):

THANK YOU!!! :)

OpenStudy (amistre64):

yep, hope it helps :)